OpenStudy (anonymous):

Help please: Rate question! A street light is at the top of a 15 ft tall pole. A woman 6 ft tall walks away from the pole with a speed of 7 ft/sec along a straight path. How fast is the tip of her shadow moving when she is 30 ft from the base of the pole?

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

We will need to identify the similar triangles. Side - Pole : Woman :: 15 ft : 6 ft Angle - Pole/Path vs. Woman/Path Side - Pole to Light at the end of the shadow vs. Woman to end of shadow. There it is! Now what?

OpenStudy (anonymous):

Yes Find the dy/dt?

OpenStudy (tkhunny):

Before you can find dy/dt, you must define 'y'. What is it?

OpenStudy (anonymous):

6x+6y=15y?

OpenStudy (anonymous):

use two tringle?

OpenStudy (tkhunny):

This is one thing you must learn. You can NOT just write down variables without defining them. What is x? What is y? No one can see into your head. You must communicate what is in there. Is y the length of the shadow?

OpenStudy (anonymous):

Ok, I draw it 

OpenStudy (tkhunny):

There you go! A clue! Now the similar triangles: \(\dfrac{6}{15} = \dfrac{y}{x+y}\) We're almost done.

OpenStudy (anonymous):

is my drawing wrong?

OpenStudy (anonymous):

6x+6y=15y?

OpenStudy (tkhunny):

Why would you ask that? Your drawing is what it is. It provides the definitions we seek. Now, go with it.

OpenStudy (tkhunny):

And now, only AFTER we see your drawing, and thus your definitions, can we understand your equation. Now what?

OpenStudy (anonymous):

Yes, beacse we are using two triangle.

hartnn (hartnn):

could you identify what 7ft/sec is ?

OpenStudy (anonymous):

speed?

hartnn (hartnn):

i know but, its dx/dt = 7ft/sec.

OpenStudy (anonymous):

No, so far 6x+6y=15y

OpenStudy (anonymous):

thats all i understand

hartnn (hartnn):

she walks away from pole with that speed, so x changes with that speed, so dx/dt =7 got this ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so theres 2 equation?

hartnn (hartnn):

now isolate y from 6x+6y=15y and differentiate it.

OpenStudy (anonymous):

6x=9y? is this how do you isolate y?

hartnn (hartnn):

yeah..then y=6x/9 now if you take derivative, with respect to 't' what do you get ?

OpenStudy (anonymous):

y=6x/9 dt/dx?

hartnn (hartnn):

no... dy/dt = 6* (dx/dt)/9 got this ?

OpenStudy (anonymous):

yes beacuse its y.... i got it..

hartnn (hartnn):

now you have dx/dt=7 and you need dy/dt can you find it ?

OpenStudy (anonymous):

d/dt=7*dy/dt?

hartnn (hartnn):

? dy/dt = 6* (dx/dt)/9 with dx/dt=7 gives you dy/dt = 6* 7/9 -_-

OpenStudy (anonymous):

Oh haha I thought it was totally saprate things! ops

hartnn (hartnn):

so, the shadow is moving with the speed of 14/3 ft/sec....